### v1

# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

Abstract
Practical Bayes filters often assume the state distribution of each time step to be Gaussian for computational tractability, resulting in the so-called Gaussian filters. When facing nonlinear systems, Gaussian filters such as extended Kalman filter (EKF) or unscented Kalman filter (UKF) typically rely on certain linearization techniques, which can introduce large estimation errors. To address this issue, this paper reconstructs the prediction and update steps of Gaussian filtering as solutions to two distinct optimization problems, whose optimal conditions are found to have analytical forms from Stein’s lemma. It is observed that the stationary point for the prediction step requires calculating the first two moments of the prior distribution, which is equivalent to that step in existing moment-matching filters. In the update step, instead of linearizing the model to approximate the stationary points, we propose an iterative approach to directly minimize the update step’s objective to avoid linearization errors. For the purpose of performing the steepest descent on the Gaussian manifold, we derive its natural gradient that leverages Fisher information matrix to adjust the gradient direction, accounting for the curvature of the parameter space. Combining this update step with moment matching in the prediction step, we introduce a new iterative filter for nonlinear systems called Natural Gradient Gaussian Approximation filter, or NANO filter for short. We prove that NANO filter locally converges to the optimal Gaussian approximation at each time step. The estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through contructing a supermartingale-like inequality across consecutive time steps. Real-world experiments demonstrate that, compared to popular Gaussian filters such as EKF, UKF, iterated EKF, and posterior linearization filter, NANO filter reduces the average root mean square error by approximately 45% while maintaining a comparable computational burden.
Index Terms:
State estimation, Bayesian filtering, Gaussian filter, natural gradient descent
I Introduction
State estimation of dynamical systems is a timely topic in fields such as astrophysics, robotics, power systems, manufacturing, and transportation. The most comprehensive framework for state estimation is Bayesian filtering, which targets the distribution of the current state given the available measurements to date. This state distribution, referred to as the posterior distribution, is often calculated iteratively through two steps, i.e., prediction and update. The prediction step uses the Chapman-Kolmogorov equation to forward-predict the state distribution using the transition probability, a conditional probability that describes the time evolution of the system state. Based on the prior distribution acquired after prediction, the update step applies Bayes’ theorem to update the prior using the measurement probability, a conditional probability that describes the relationship between noisy measurements and the true state [1].
For linear Gaussian systems, directly applying Bayesian filtering results in the well-known Kalman filter (KF) [2], which analytically computes the Gaussian posterior through recursive updates of its mean and covariance. This analytical form relies on the closure property of Gaussian under linear transformations and the conjugate property of Gaussian under conditioning. Unfortunately, such an elegant structure of KF does not exist for nonlinear systems, as the closure and conjugate properties only hold in the linear Gaussian case.
Therefore, if the system is nonlinear or non-Gaussian, finding the exact solution of Bayesian filtering is often unattainable. In this case, designing an appropriate approximation of state distribution becomes a key step in Bayesian filtering. To this effect, the particle filter (PF) approximates the probability density function with a set of discrete particles, each representing a possible state of the system. These particles, when weighted and summed, form a discrete approximation of the continuous probability density function. Although PF can provide asymptotically optimal approximations, it requires a large number of particles, which results in significant computational demands and limits its application in practical systems [3, 4]. In contrast to approximating the posterior distribution with samples, an alternative choice is to approximate it as Gaussian distribution in each time step, which leads to the Gaussian filter family. Compared to PF, Gaussian filters offer higher computational efficiency and have become by far the most popular family to date [5].
It turns out that the design philosophy of Gaussian filters generally consists of two steps: (i) approximate the nonlinear models to linear forms with additive Gaussian noise, and (ii) perform KF based on this linear Gaussian model. Within this framework, the differences between Gaussian filters are primarily attributed to different linearization techniques. The earliest technique involves directly using the Taylor series approximation to linearize the nonlinear function. Typically, the first-order expansion is employed to avoid the occurrence of tensors in high-order expansions, exemplified by the well-known extended Kalman filter (EKF) [6, 7]. Originally developed by NASA for navigation tasks, EKF linearizes the nonlinear function around the state estimate and performs KF using the resulting affine-form system equation. Building on this foundation, iterated extended Kalman filter (IEKF) [8] repeatedly performs linearization at each updated approximation of the posterior mean instead of the prior mean as in EKF. In fact, the iterative process of this method is a Gauss-Newton iteration that essentially linearizes the model around the maximum a posteriori (MAP) estimate of the state [9].
A clear drawback of directly linearizing the state space model is its inability to capture the second moment, namely the covariance, after a nonlinear transformation, since Taylor-series linearization only utilizes the mean. To address this issue, a fundamentally different approach is to directly match the transformed mean and covariance, known as moment-matching. Essentially, moment-matching can be viewed as statistical linear regression (SLR), a linearization technique that obtains optimal affine representation of the nonlinear system with parameters minimizing the expected regression loss [10]. This so-called SLR method requires computing several integrals over Gaussian distributions. In particular, the use of established numerical integration methods such as unscented transform, Gauss–Hermite integration, and spherical cubature integration underpin the design of unscented Kalman filter (UKF) [11], Gauss–Hermite KF [12], and cubature KF [13]. In addition, a new algorithm called posterior linearization filter (PLF) is proposed to perform SLR at the posterior rather than the prior during Bayesian updates to improve filtering accuracy [10].
This framework of first linearizing the model and then performing KF is termed enabling approximation [10]. Although it has been widely applied since the 1960s and may seem like a natural choice, one critical question cannot be overlooked: Is this enabling approximation framework sufficient to find the optimal Gaussian approximation of Bayesian filtering? Unfortunately, no existing works discuss this issue. In this paper, we argue that applying this framework in the update step may not yield an exact solution for nonlinear Gaussian filtering. This contrasts with the prediction step, where performing enabling approximation using moment-matching filters [11, 12, 13] is a proper choice. To this effect, we propose a new method called Natural grAdient GaussiaN apprOximation (NANO) filter that applies natural gradient in the update step to find the exact solution of Gaussian approximation. Specifically, the contributions of this paper are summarized as follows:
  • We interpret the prediction and update steps of Bayesian filtering as solutions to two distinct optimization problems. This new perspective allows us to define optimal Gaussian approximation and identify its corresponding extremum conditions. Leveraging the Stein’s lemma, we derive that the stationary point for prediction step has an explicit form, involving the calculation of the first two moments of prior distribution. This analytical form is implementable via moment-matching filters [11, 12, 13]. In contrast, the stationary point for update step is characterized by two interdependent equations, which generally has no analytical root. A special case occurs in the linear Gaussian systems, where these two equations decouple and become linear, resulting in analytical solutions for KF. For nonlinear systems, those enabling approximation-based Gaussian filters need to perform certain linearization technique, which inevitably introduce large estimation errors.
  • To address linearization errors in the update step, we derive natural gradient iteration to minimize the optimization cost. By leveraging the Fisher information matrix that captures the curvature of the parameter space, the gradient direction is adjusted to perform the steepest descent on the Gaussian manifold. By combining this optimization procedure in the update step with moment matching in the prediction step, we develop a new iterative filter for nonlinear systems, namely NANO filter. We demonstrate that KF is equivalent to a single iteration of NANO filter for linear Gaussian systems, providing a fresh understanding of Kalman filtering.
  • We prove that the NANO filter locally converges to the optimal Gaussian approximation at each time step, with accuracy up to a second-order remainder in the Taylor expansion. Additionally, the estimation error is proven exponentially bounded for nearly linear measurement equation and low noise levels through contructing a supermartingale-like inequality across consecutive time steps. We further show that NANO filter can naturally extend the Bayesian posterior to Gibbs posterior, enabling the use of more flexible loss functions to enhance robustness against outliers in measurement data. On its basis, three robust variants of NANO filter are introduced, each employing different robust loss function designs. Simulations and real-world experiments demonstrate that NANO filter and its robust variants significantly outperforms popular Gaussian filters, such as EKF, UKF, IEKF, and PLF for nonlinear systems.
The remainder of this paper is structured as follows: Section II formulates the problem, and Section III explores the optimal Gaussian approximation. Section IV introduces the NANO filter algorithm, followed by theoretical analysis in Section V. Section VI proposes the robust variants of NANO filter. Section VII provides a discussion on the proposed algorithm. Simulations and real-world experiments are presented in Section VIII.
Notation: All vectors are considered as column vectors. The symbol $D_{\mathrm{KL}}(p\|q)=\mathbb{E}_{p}\left\{\log{\frac{p}{q}}\right\}$ denotes the KL divergence between two probability distributions $p$ and $q$. Unless otherwise specified, $\|x\|$ refers to the $\ell_{2}$-norm of the vector $x$, defined as $\|x\|=\sqrt{x^{\top}x}$. The notation $\mathcal{N}(x;\mu,\Sigma)$ represents the Gaussian probability density function for the variable $x$ with mean $\mu$ and covariance matrix $\Sigma$. For simplicity, this may be abbreviated as $\mathcal{N}(\mu,\Sigma)$. Furthermore, $\mathbb{E}\{x\}$, $\mathbb{D}\{x\}$, and $\mathrm{Cov}(x,y)$ denote the expectation, variance, and covariance, respectively. The notation $A\leq B$ ($A<B$) for symmetric matrices $A$ and $B$ indicates that $B-A$ is positive (semi-)definite. The symbol $\otimes$ denotes the Kronecker product.
II Problem Statement
Consider the following nonlinear discrete-time stochastic system:
$x_{t+1}=f(x_{t})+\xi_{t},$ (1)
$y_{t}=g(x_{t})+\zeta_{t},$
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $y_{t}\in\mathbb{R}^{m}$ is the noisy measurement. The function ${f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is referred to as the transition function, while ${g}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is called measurement function; ${\xi}_{t}$ denotes process noise, and ${\zeta}_{t}$ represents measurement noise. Typically, the initial state $x_{0}$, the process noise $\{\xi_{t}\}$, and the measurement noise $\{\zeta_{t}\}$ are all mutually independent, with the process and measurement noises being independent and identically distributed across time. This state-space model description in (1) can be transformed into a hidden Markov model:
$x_{0}$ $\sim p(x_{0}),$ (2)
$x_{t}$ $\sim p(x_{t}|x_{t-1}),$
$y_{t}$ $\sim p(y_{t}|x_{t}).$
Here, $p(x_{t}|x_{t-1})$ and $p(y_{t}|x_{t})$ represent the transition and output probabilities respectively while $p(x_{0})$ denotes the initial state distribution. In essence, (2) and (1) are different representations of the same system model. For example, consider the transition model $x_{t}=Ax_{t-1}+\xi_{t-1}$, where $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, with $Q$ denoting the covariance matrix of the process noise. This can be equivalently described by the transition probability $p(x_{t}|x_{t-1})=\mathcal{N}(x_{t};Ax_{t-1},Q)$.
The objective of state estimation is to recover the system state $x_{t}$ from noisy measurements $y_{t}$. Typically, finding the optimal estimate involves two key steps: calculating the posterior distribution $p(x_{t}|y_{1:t})$, and determining the optimal estimate $\hat{x}_{t|t}$ using the posterior distribution. A principled framework for posterior distribution calculation is Bayesian filtering, which computes $p(x_{t}|y_{1:t})$ recursively through two steps:
$p(x_{t}|y_{1:t-1})$ $=\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1},$ (3a)
$p(x_{t}|y_{1:t})$ $=\frac{p(y_{t}|x_{t})p(x_{t}|y_{1:t-1})}{\int p(y_{t}|x_{t})p(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}}.$ (3b)
Here, (3a) is called prediction step while (3b) is called update step. The former, known as the Chapman-Kolmogorov equation, utilizes the transition probability $p(x_{t}|x_{t-1})$ to predict the prior distribution $p(x_{t}|y_{1:t-1})$. Based on this prior, the update step leverages the Bayes formula to calculate the posterior distribution $p(x_{t}|y_{1:t})$, where the output probability serves as likelihood function. After calculating $p(x_{t}|y_{1:t})$, common estimation criteria like the minimum mean square error or maximum a posteriori estimation can be used to determine the optimal estimate.
As discussed in Section I, neither (3a) nor (3b) can be calculated analytically when the transition or output probabilities are non-linear or non-Gaussian, requiring the approximation of state distributions in practice. Given the high computational burden of discrete approximations like PF, Gaussian approximations are widely adopted in industrial applications. More specifically, in Gaussian filters, both the prior and posterior distributions are approximated by Gaussian distributions:
$p(x_{t}|y_{1:t-1})$ $\approx\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1}),$ (4)
$p(x_{t}|y_{1:t})$ $\approx\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t}),$
which is usually achieved through enabling approximation framework. In particular, this framework first approximates the transition or output probabilities with linear Gaussian models and then performs the well-known KF. Due to the closure property of Gaussians under linear transformations and the conjugate property of Gaussians under conditioning, the resulting prior and posterior distributions are naturally approximated as Gaussian. The linearization techniques in existing Gaussian filters can be categorized into two types: Taylor series expansion and stochastic linear regression. As summarized in Table I, the former uses the Jacobian matrix $g^{\prime}(\bar{x})$ to provide local affine approximations of system models at point $\bar{x}$, while the latter minimizes the expectation of the square of the residual $y-Ax-b$ to find optimal linear parameters.
TABLE I: Linearization techniques for existing Gaussian filters
Linearization Technique Basic Principle $\mathcal{N}(y;g(x),\Sigma)\approx\mathcal{N}(y;Ax+b,\Lambda)$ Representative Algorithms
Taylor Series Expansion $g(x)=g(\bar{x})+g^{\prime}(\bar{x})(x-\bar{x})$Downarrow $A=g^{\prime}(\bar{x}),b=g(\bar{x})-g^{\prime}(\bar{x})\bar{x},\Lambda=\Sigma$ EKF[6], IEKF[8]
Statistical Linear Regression $\operatorname*{arg\,min}_{A,b}\mathbb{E}_{x}\left\{(y-Ax-b)(y-Ax-b)^{\top}\right\}$Downarrow $A=\bm{\mathrm{Cov}}(x,y)^{\top}\mathbb{D}(x),b=\mathbb{E}\left\{y\right\}-A\mathbb{E}\left\{x\right\},\Lambda=\mathbb{D}\left\{y\right\}-A\mathbb{D}\left\{x\right\}A^{\top}$ UKF [11], Gauss–Hermite KF [12], cubature KF [13], and PLF[14]
While these techniques offer practical methods for Gaussian approximation, it still remains unclear whether this enabling approximation framework truly achieve an optimal Gaussian approximation for Bayesian filtering. Unfortunately, there is even no clear way to formally judge whether a Gaussian approximation is optimal. This paper aims to establish a framework for defining and identifying the optimal Gaussian approximation for Bayesian filtering. Specifically, we seek to address the following three questions:
  1. Q1:
    What defines the optimal Gaussian approximation in Bayesian filtering, and what conditions must it meet?
  2. Q2:
    How can we design a filter that effectively achieves this optimal approximation?
  3. Q3:
    What theoretical guarantees can we provide for the algorithm’s convergence and the boundedness of the estimation error?
III Optimal Gaussian Approximation for Bayesian Filtering
In this section, we will address Q1 by exploring the necessary conditions for determining the optimal Gaussian approximation for Bayesian filtering.
III-A Optimization Viewpoint of Bayesian Filtering
Inspired by the optimization-centric view on Bayes’s rule [15], we show that the prior and posterior distributions can be interpreted as solutions of two variational problems, as shown in the subsequent proposition.
Proposition 1 (Variational Problems for Bayesian filtering).
The prior distribution can be regraded as the maxmizer of an variational problem:
$p(x_{t}|y_{1:t-1})=\operatorname*{arg\,max}_{q(x_{t})}\mathbb{E}_{\begin{subarray}{c}p(x_{t-1}|y_{1:t-1})\\ p(x_{t}|x_{t-1})\end{subarray}}\left\{\log{q(x_{t})}\right\}.$ (5)
Similarly, the posterior distribution can be regarded as the minimizer of a functional:
$p(x_{t}|y_{1:t})=$ $\operatorname*{arg\,min}_{q(x_{t})}\big{\{}\mathbb{E}_{q(x_{t})}\left\{-\log{p(y_{t}|{x}_{t})}\right\}$ (6)
$+D_{\mathrm{KL}}\left(q(x_{t})||p(x_{t}|y_{1:t-1})\right)\big{\}}.$
Note that in both (5) and (6), $q:\mathbb{R}^{n}\to\mathbb{R}$ represents the candidate density function. Besides, we use the notation $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}$ to represent the expectation of $f(x,y)$ with respect to both distributions $p(x)$ and $p(y)$, i.e., $\mathbb{E}_{\begin{subarray}{c}p(x)\\ p(y)\end{subarray}}\left\{f(x,y)\right\}:=\mathbb{E}_{p(x)}\mathbb{E}_{p(y)}\left\{f(x,y)\right\}$.
Proof.
By recognizing that the focus is solely on the extremizer and not on the objective value itself, it follows that for any constant $Z>0$, the right-hand side of (5) can be rewritten as
$\operatorname*{arg\,min}_{q(x_{t})}\left\{\mathbb{E}_{\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1}}\left\{\log{\frac{Z}{q(x_{t})}}\right\}\right\}.$ (7)
If we choose $Z=\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1}$ and recall that the Kullback-Leibler (KL) divergence reaches its minimum value uniquely when its arguments are identical, we find that the solution to (7) is indeed
$\int p(x_{t}|x_{t-1})p(x_{t-1}|y_{1:t-1})\,\mathrm{d}x_{t-1},$
which corresponds exactly to the prior distribution defined in (3a). Similarly, the right-hand side of (6) is equal to
$\operatorname*{arg\,min}_{q(x_{t})}\left\{\mathbb{E}_{q(x_{t})}\left\{\log{\frac{q(x_{t})}{p(y_{t}|x_{t})p(x_{t}|y_{1:t-1})/Z}}\right\}\right\}.$ (8)
If we choose $Z=\int p(y_{t}|x_{t})p(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}$, the solution to (8) is
$\frac{p(y_{t}|x_{t})p(x_{t}|y_{1:t-1})}{\int p(y_{t}|x_{t}){p}(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}},$
which is precisely the posterior distribution defined in (3b). ∎
This proposition shows that the prior distribution $p(x_{t}|y_{1:t-1})$ can be viewed as the solution to a variational problem that maximizes the expected logarithm of a candidate density $q(x_{t})$ over the joint distribution of the previous state and the transition probability. This reflects the idea that the prior is derived by considering all possible transitions from the previous state, and selecting the distribution that maximizes the expected log-density under these transitions. Similarly, the posterior distribution $p(x_{t}|y_{1:t})$ is the solution to a variational problem that minimizes a cost combining the expected negative log-likelihood of the measurement model, $-\log{p(y_{t}|{x}_{t})}$, with the KL divergence between the candidate density and the prior distribution $D_{\mathrm{KL}}(q(x_{t})||p(x_{t}|y_{1:t-1}))$. This captures the Bayesian update process, where the posterior distribution adjusts the prior distribution based on new measurements to balance prior knowledge with new information.
In both (5) and (6), there are no constraints on the candidate distribution, meaning we seek an optimal candidate distribution over the entire probability space. However, such variational problems generally lack analytical solutions. A tractable approach to solve this problem is to restrict the candidate distribution to a parameterizable family of distributions, with the Gaussian distribution family being the most commonly used [16]. Therefore, by utilizing Gaussian approximations in (4), the two variational problems (5) and (6) that depict Bayesian filtering can be reduced to two optimization problems:
$\hat{x}_{t|t-1},P_{t|t-1}=$ $\operatorname*{arg\,max}_{\hat{x}_{t},P_{t}}L(\hat{x}_{t},P_{t}),$
$L(\hat{x}_{t},P_{t})=$ $\mathbb{E}_{\begin{subarray}{c}\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})\\ p(x_{t}|x_{t-1})\end{subarray}}\left\{\log{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\right\},$ (9a)
$\hat{x}_{t|t},P_{t|t}=$ $\operatorname*{arg\,min}_{\hat{x}_{t},P_{t}}J(\hat{x}_{t},P_{t}),$
$J(\hat{x}_{t},P_{t})=$ $D_{\mathrm{KL}}\left(\mathcal{N}(x_{t};\hat{x}_{t},P_{t})||\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})\right)$
$-\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\big{\{}\log{p(y_{t}|{x}_{t})}\big{\}}.$ (9b)
Here, $L(\hat{x}_{t},P_{t})$ is called prediction cost while $J(\hat{x}_{t},P_{t})$ is called update cost. Compared with (5) and (6), (9a) and (9b) transform the generally unsolvable variational problems to optimization problems for Gaussian parameters. This allows us to find the optimal Gaussian approximation for Bayesian filtering by studying the optimality conditions of (9).
III-B Structure of Stationary Points of $L(\hat{x}_{t},P_{t})$
To establish the optimality conditions for problems (9a) and (9b), we examine the stationary points of $L(\hat{x}_{t},P_{t})$ and $J(\hat{x}_{t},P_{t})$. Interestingly, the extreme condition of the former reduces to a straightforward moment-matching equation, while that of the latter results in two mutually coupled implicit equations, whose roots are generally intractable. The following lemma helps elucidate the structure of $L(\hat{x}_{t},P_{t})$:
Lemma 1 (Stationary Points for Maximum Gaussian Likelihood).
For probability density function $p(x)$, the stationary points of a maximum expected Gaussian likelihood problem
$\mu^{*},\Sigma^{*}=\operatorname*{arg\,max}_{\mu,\Sigma}\mathbb{E}_{p(x)}\left\{\log\mathcal{N}(x;\mu,\Sigma)\right\},$ (10)
can be written as
$\mu^{*}$ $=\mathbb{E}_{p(x)}\left\{x\right\},$ (11a)
$\Sigma^{*}$ $=\mathbb{E}_{p(x)}\left\{(x-\mu^{*})(x-\mu^{*})^{\top}\right\}$ (11b)
$=\mathbb{E}_{p(x)}\left\{xx^{\top}\right\}-\mu^{*}{\mu^{*}}^{\top}.$ (11c)
The proof of this lemma can be found in Appendix -A. This lemma states that the stationary points for maximizing the expected Gaussian likelihood are achieved when the Gaussian distribution matches the mean and variance of the given distribution. Therefore, as a special case of (10), solving (9a) requires to match the first and second order moment:
$\hat{x}_{t|t-1}$ $=\mathbb{E}_{\begin{subarray}{c}\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})\\ p(x_{t}|x_{t-1})\end{subarray}}\left\{x_{t}\right\},$ (12)
$P_{t|t-1}$ $=\mathbb{E}_{\begin{subarray}{c}\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})\\ p(x_{t}|x_{t-1})\end{subarray}}\left\{x_{t}x_{t}^{\top}\right\}-\hat{x}_{t|t-1}\hat{x}_{t|t-1}^{\top}.$
For the system obeys the transition model in (1) with $\xi_{t}$ being the zero mean process noise, (12) can be further expressed as
$\hat{x}_{t|t-1}=$ $\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{f(x_{t-1})\right\},$ (13)
$P_{t|t-1}=$ $\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{f(x_{t-1})f^{\top}(x_{t-1})\right\}$
$+\mathbb{D}\left\{\xi_{t-1}\right\}-\hat{x}_{t|t-1}\hat{x}_{t|t-1}^{\top}.$
By leveraging variable substitution and the independence between process noise and state, we find that the optimal solution in (13) essentially requires finding the expectation of a nonlinear function with respect to a Gaussian distribution. The subsequent example explains this momoment-matching operation in terms of the canonical KF:
Example 1 (Prediction step of Kalman filter).
For linear Gaussian systems $x_{t}=Ax_{t-1}+\xi_{t-1}$ where the process noise satisfies $\xi_{t-1}\sim\mathcal{N}(\xi_{t-1};0,Q)$, if the posterior distribution at time $t-1$ is given by $x_{t-1}\sim\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})$, the prediction step using (13) yields the following stationary points:
$\hat{x}_{t|t-1}=$ $\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{Ax_{t-1}\right\}$
$=$ $A\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{x_{t-1}\right\}$
$=$ $A\hat{x}_{t-1|t-1},$
$P_{t|t-1}=$ $\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{Ax_{t-1}x_{t-1}^{\top}A^{\top}\right\}$
$+\mathbb{D}\left\{\xi_{t-1}\right\}-A\hat{x}_{t-1|t-1}\hat{x}_{t-1|t-1}^{\top}A^{\top},$
$=$ $AP_{t-1|t-1}A^{\top}+Q.$
Note that the prior covariance matrix satisfies $P_{t-1|t-1}=\mathbb{E}_{\mathcal{N}(x_{t-1};\hat{x}_{t-1|t-1},P_{t-1|t-1})}\left\{x_{t-1}x_{t-1}^{\top}\right\}-\hat{x}_{t-1|t-1}\hat{x}_{t-1|t-1}^{\top}$.
In linear systems as shown in Example 1, the expectation operator in (13) is allowed to be interchanged with the affine function. However, for nonlinear functions $f$, computing the expectation requires numerical methods. As discussed in Section I, moment-matching KF methods, such as UKF, GHKF, and CKF, approximate this expectation using techniques like the unscented transform, Gauss–Hermite quadrature, and spherical cubature. These methods provide optimal Gaussian approximations for the prediction step by numerically solving the expectation.
III-C Structure of Stationary Points of $J(\hat{x}_{t},P_{t})$
Compared to the simple form of the stationary points of $L(\hat{x}_{t},P_{t})$, the stationary points of $J(\hat{x}_{t},P_{t})$ are relatively more complicated. For simplicity in notation, we define the negation of the log-likelihood as $\ell(x_{t},y_{t})$, where $\ell(x_{t},y_{t})=-\log{p(y_{t}|x_{t})}$, and refer to $\ell(x_{t},y_{t})$ as the measurement-dependent loss. Using the analytical form of the KL divergence for two Gaussian distributions, the update cost can be formulated as
$J(\hat{x}_{t},P_{t})$ (14)
$=$ $\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\left\{\ell(x_{t},y_{t})\right\}$
$+D_{\text{KL}}\left(\mathcal{N}(x_{t};\hat{x}_{t},P_{t})||\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})\right)$
$=$ $\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\left\{\ell(x_{t},y_{t})\right\}$
$+\frac{1}{2}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)^{\mathrm{T}}P_{t|t-1}^{-1}\left(\hat{x}_{t|t-1}-\hat{x}_{t}\right)$
$+\frac{1}{2}\mathrm{tr}\left(P_{t|t-1}^{-1}P_{t}\right)-\frac{1}{2}\log\frac{\left|P_{t}\right|}{\left|P_{t|t-1}\right|}-\frac{1}{2}n.$
To find the stationary points of $J(\hat{x}_{t},P_{t})$, we need to calculate the partial derivatives with respect to $\hat{x}_{t}$ and $P_{t}$. The following lemma is helpful in simplifying these partial derivative calculations:
Lemma 2 (Gradient of expectation under Gaussian distribution).
Assuming that $f:\mathbb{R}^{n}\to\mathbb{R}$ is twice differentiable, we have the following results:
(i). The gradient of $\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}$ w.r.t. the mean $\mu$ satisfies
$\frac{\partial}{\partial\mu}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}$ $=\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{\frac{\partial}{\partial x}f(x)\right\}$ (15)
$=\Sigma^{-1}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{(x-\mu)f(x)\right\}.$
(ii). The Hessian matrix of $\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}$ w.r.t. the mean $\mu$ satisfies
$\frac{\partial^{2}}{\partial\mu^{2}}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}$ (16)
$=$ $\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{\frac{\partial^{2}}{\partial x^{2}}f(x)\right\}$
$=$ $-2\Sigma^{-1}\left(\frac{\partial}{\partial\Sigma^{-1}}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}\right)\Sigma^{-1}$
$=$ $\Sigma^{-1}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{(x-\mu)(x-\mu)^{\top}f(x)\right\}$
$-\Sigma^{-1}\mathbb{E}_{\mathcal{N}(x;\mu,\Sigma)}\left\{f(x)\right\}.$
Proof.
The results presented are all related to Stein’s lemma [17]. Result (i) is known as Bonnet’s Theorem [18] and result (ii) is referred to as Price’s Theorem [19]. Detailed discussions about these results can be found in this technical report [20]. ∎
Before studying the structure of the stationary points, we have the following assumption:
Assumption 1.
The measurement-dependent loss function $\ell(x_{t},y_{t})$ is twice differentiable with respect to the state $x_{t}$.
Based on Assumption 1 and the result of Lemma 2, we have the partial derivative of $J(\hat{x}_{t},P_{t})$ with respect to $\hat{x}_{t}$ and $P_{t}^{-1}$:
$\frac{\partial J(\hat{x}_{t},P_{t})}{\partial\hat{x}_{t}}$
$=$ $\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\left\{\frac{\partial\ell(x_{t},y_{t})}{\partial x_{t}}\right\}+P_{t|t-1}^{-1}(\hat{x}_{t}-\hat{x}_{t|t-1}),$ (17a)
$\frac{\partial J(\hat{x}_{t},P_{t})}{\partial P_{t}^{-1}}$
$=$ $-\frac{1}{2}P_{t}\cdot\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial x_{t}^{2}}\right\}\cdot P_{t}$
$-\frac{1}{2}P_{t}P^{-1}_{t|t-1}P_{t}+\frac{1}{2}P_{t}.$ (17b)
To find extrema, we could attempt to set the first derivatives to zero
$\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{\hat{x}_{t|t}}}=\frac{\partial{J}(\hat{x}_{t|t},P_{t|t})}{\partial{P_{t|t}}}=0,$
we have
$\hat{x}_{t|t}$ $=\hat{x}_{t|t-1}-P_{t|t-1}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\left\{\frac{\partial\ell(x_{t},y_{t})}{\partial x_{t}}\right\},$ (18a)
$P_{t|t}^{-1}$ $=P_{t|t-1}^{-1}+\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial x_{t}^{2}}\right\}.$ (18b)
As shown in (18), the first-order condition is generally not possible to isolate for $\hat{x}_{t|t}$ and $P_{t|t}$. An exception for it is the well-known KF, where the expectation in (18) is a constant value, as shown in the subsequent example:
Example 2 (Update step of Kalman filter).
For linear Gaussian systems with output probability $p(y_{t}|x_{t})=\mathcal{N}(y_{t};Cx_{t},R)$, the optimal Gaussian approximation of the posterior mean in (18a) can be written as
$\hat{x}_{t|t}$ (19)
$=$ $\hat{x}_{t|t-1}+P_{t|t-1}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\left\{\frac{\partial}{\partial{x_{t}}}\log\left\{\mathcal{N}(y_{t};Cx_{t},R)\right\}\right\}$
$=$ $\hat{x}_{t|t-1}-P_{t|t-1}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\left\{\frac{\partial}{\partial{x_{t}}}\left\{\frac{1}{2}\|y_{t}-Cx_{t}\|^{2}_{R^{-1}}\right\}\right\}$
$=$ $\hat{x}_{t|t-1}+P_{t|t-1}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\left\{C^{\top}R^{-1}(y_{t}-Cx_{t})\right\}$
$=$ $\hat{x}_{t|t-1}+P_{t|t-1}C^{\top}R^{-1}(y_{t}-C\hat{x}_{t|t}).$
From (19), we can obtain
$\hat{x}_{t|t}=\hat{x}_{t|t-1}+K_{t}(y_{t}-C\hat{x}_{t|t-1}),$ (20)
with $K_{t}$ being the gain matrix of KF defined as $K_{t}\triangleq P_{t|t-1}C^{\top}(CP_{t|t-1}C^{\top}+R)^{-1}$. Similarly, the optimal Gaussian approximation of the posterior covariance (18b) can be written as
$P_{t|t}$ (21)
$=$ $\Big{(}P_{t|t-1}^{-1}-\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\big{\{}\frac{\partial^{2}}{\partial{x_{t}^{2}}}\log\left\{\mathcal{N}(y_{t};Cx_{t},R)\right\}\big{\}}\Big{)}^{-1}$
$=$ $\left(P_{t|t-1}^{-1}+C^{\top}R^{-1}C\right)^{-1}$
$=$ $P_{t|t-1}-P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1}CP_{t|t-1}.$
Note that (20) and (21) consist of the update step of the canonical KF.
As shown in Example 2, for the special case of a linear Gaussian system, the extreme conditions are specified by two decoupled equations whose analytical solutions correspond to the analytical form of the KF update. Specifically, the expectation on the right-hand side of (18a) depends solely on the posterior mean, while the expectation in (18b) is determined entirely by the system, meaning it is unrelated to either the mean or covariance. However, for a general nonlinear or non-Gaussian system, the expectations in (18) depend on both the posterior mean and covariance, rendering the stationary points analytically intractable.
Remark 1.
As discussed in Section II, existing Gaussian filters rely on approximation techniques that solve (18) by linearizing the measurement model and then performing the KF updates as (20) and (21). We contend that this type of approximation technique inevitably introduces linearization errors. For example, the Taylor series expansion technique unavoidably results in higher-order error terms. Moreover, these approximation techniques are applicable only for Gaussian noises [10]. Therefore, there is an urgent need to develop a new method that can directly solve the Gaussian approximation for Bayesian filtering update to avoid linearization errors.
IV Natural Gradient Gasussian Approximation
In the previous section, we observed that existing moment-matching KF methods already provide approximate solutions for the optimal Gaussian approximation in the prediction step, but the update step remains inadequately addressed in existing Gaussian filters. In this section, we aim to tackle Q2. Specifically, we will design an algorithm to solve (9b).
By examining the structure of the extreme conditons defined by (18), we find that it is challenging to directly obtain an analytical form of the stationary point because it is typically impossible to isolate the updates for the mean and covariance in (18). Therefore, a more practical approach to finding the optimal solution is to directly minimize the update cost $J(\hat{x}_{t},P_{t})$. To find the steepest descent in optimizing the parameters of Gaussian distributions [21, 22], we derive an natural gradient iteration for finding the optimal Gaussian approximation.
For simplicity, we stack the Gaussian parameters into a single column vector $v$ and calculate the derivative with respect to it:
$v=\begin{bmatrix}\hat{x}_{t}\\ \operatorname*{\mathrm{vec}}(P_{t}^{-1})\end{bmatrix},\;\frac{\partial}{\partial{v}}J(\hat{x}_{t},P_{t})$ $=\begin{bmatrix}\frac{\partial}{\partial\hat{x}_{t}}J(\hat{x}_{t},P_{t})\\ \operatorname*{\mathrm{vec}}\left(\frac{\partial}{\partial P_{t}^{-1}}J(\hat{x}_{t},P_{t})\right)\end{bmatrix}.$ (22)
Here, we consider the inverse of the covariance matrix instead of its original form. This consideration is inspired by the structure of the information filter [23, 24], an equivalent form of the KF, where the inverse of the covariance matrix is employed instead of the covariance matrix itself. This is because the inverse can potentially simplify the mathematical expression of the update step in Bayesian filtering [23, 24]. Additionally, to easily represent the iteration, we define
$\delta v\triangleq\begin{bmatrix}\delta\hat{x}_{t}\\ \operatorname*{\mathrm{vec}}(\delta P_{t}^{-1})\end{bmatrix}=\begin{bmatrix}\hat{x}_{t}^{(i+1)}-\hat{x}_{t}^{(i)}\\ \mathrm{vec}\left(\left(P_{t}^{-1}\right)^{(i+1)}-\left(P_{t}^{-1}\right)^{(i)}\right)\end{bmatrix},$
where $i$ is the iteration index. Under this notation, the natural gradient parameter update can be defined as
$\delta v=-\left[\mathcal{F}_{v}^{-1}\frac{\partial}{\partial v}J(\hat{x}_{t},P_{t})\right]_{v=v^{(i)}},$ (23)
where $\mathcal{F}_{v}$ is the fisher information matrix associated with the Gaussian distribution $\mathcal{N}(x_{t};\hat{x}_{t},P_{t})$ and $v^{(i)}$ represents the value of $v$ in the $i$-th iteration. The next proposition provides the formulation of Fisher information matrix:
Proposition 2.
The inverse of the fisher information matrix $\mathcal{F}^{-1}_{v}$ associated with $\mathcal{N}(x_{t};\hat{x}_{t},P_{t})$ is
$\mathcal{F}^{-1}_{v}=\begin{bmatrix}P_{t}&0\\ 0&2(P_{t}^{-1}\otimes P_{t}^{-1})\end{bmatrix},$ (24)
where $\otimes$ is the kronecker product.
The proof of this proposition can be found in Section 2.2 of [25]. Combining (22) and (24) with (23), we have
$\delta\hat{x}_{t}$ $=-\left[P_{t}\frac{\partial}{\partial\hat{x}_{t}}J(\hat{x}_{t},P_{t})\right]_{v=v^{(i)}},$
$\operatorname*{\mathrm{vec}}\left(\delta P_{t}^{-1}\right)$ $=-2\left[\left(P_{t}^{-1}\otimes P_{t}^{-1}\right)\operatorname*{\mathrm{vec}}\left(\frac{\partial}{\partial P_{t}^{-1}}J(\hat{x}_{t},P_{t})\right)\right]_{v=v^{(i)}}.$
Transforming this into matrix form, we can derive the following iterative updates:
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $\left(P_{t}^{-1}\right)^{(i)}$
$-2\left(P_{t}^{-1}\right)^{(i)}\left.\frac{\partial}{\partial P_{t}^{-1}}J(\hat{x}_{t},P_{t})\right|_{v^{(i)}}\left(P_{t}^{-1}\right)^{(i)},$
$\hat{x}^{(i+1)}_{t}=$ $\hat{x}^{(i)}_{t}-P_{t}^{(i+1)}\left.\frac{\partial}{\partial\hat{x}_{t}}J(\hat{x}_{t},P_{t})\right|_{v^{(i)}}.$
Combining the iterative updates (25) with the formulation of partial derivative, we have
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $P_{t|t-1}^{-1}+\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial x_{t}^{2}}\right\},$ (26)
$\hat{x}^{(i+1)}_{t}=$ $\hat{x}^{(i)}_{t}-P_{t}^{(i+1)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\frac{\partial\ell(x_{t},y_{t})}{\partial x_{t}}\right\}$
$-P_{t}^{(i+1)}P_{t|t-1}^{-1}(\hat{x}_{t}^{(i)}-\hat{x}_{t|t-1}).$
One practical issue when performing (26) is that the derivatives of $\ell(x_{t},y_{t})$ can be hard to compute. To avoid the need to compute derivatives of the measurement-dependent loss, we can once again apply Lemma 2 to acquire the derivative-free formulation. By applying (15) and (16), we have
$\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\frac{\partial\ell(x_{t},y_{t})}{\partial x_{t}}\right\}$ (27)
$=$ $\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\left(x_{t}-x_{t}^{(i)}\right)\ell(x_{t},y_{t})\right\},$
$\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial x_{t}^{2}}\right\}$
$=$ $\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\big{\{}(x_{t}-\hat{x}_{t}^{(i)})(x_{t}-\hat{x}_{t}^{(i)})^{\top}\ell(x_{t},y_{t})\big{\}}$
$-\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\ell(x_{t},y_{t})\right\}.$
With the result in (27), we have the derivative-free update scheme shown in (28).
$\left(P_{t}^{-1}\right)^{(i+1)}=$ $P_{t|t-1}^{-1}+\left(P_{t}^{-1}\right)^{(i)}\cdot\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\left(x_{t}-\hat{x}_{t}^{(i)}\right)\left(x_{t}-\hat{x}_{t}^{(i)}\right)^{\top}\ell(x_{t},y_{t})\right\}-\left(P_{t}^{-1}\right)^{(i)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\ell(x_{t},y_{t})\right\},$ (28)
$\hat{x}^{(i+1)}_{t}=$ $\hat{x}^{(i)}_{t}-P_{t}^{(i+1)}\cdot\left(P_{t}^{-1}\right)^{(i)}\cdot\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\left(x_{t}-x_{t}^{(i)}\right)\ell(x_{t},y_{t})\right\}-P_{t}^{(i+1)}P_{t|t-1}^{-1}(\hat{x}_{t}^{(i)}-\hat{x}_{t|t-1}).$
This update scheme is still practically intractable for two reasons. First, the expectations generally do not have analytical forms. Second, the update scheme in (28) generally cannot guarantee that the covariance matrix will be positive definite.
To address the first issue, we could use well-established numerical integration methods, such as the unscented transform [11], Gauss–Hermite integration [26], or spherical cubature integration [13] to approximate the expectation calculations. For the second issue, one possible and efficient solution is to provide a sufficiently good initialization. For example, we could solve the MAP estimation problem:
$\hat{x}_{t|t}^{\mathrm{MAP}}$ $=\operatorname*{arg\,max}_{x_{t}}\left\{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})\cdot\exp\{\ell(x_{t},y_{t})\}\right\}$ (29)
$=\operatorname*{arg\,max}_{x_{t}}\left\{\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}+\ell(x_{t},y_{t})\right\},$
and use Laplace’s approximation [27] to construct the initial mean and covariance in (28), as shown in the subsequent equation:
$\hat{x}_{t}^{(0)}=$ $\hat{x}_{t|t}^{\mathrm{MAP}},$ (30)
$\left(P_{t}^{-1}\right)^{(0)}=$ $\left.\frac{\partial^{2}\left\{-\log{\mathcal{N}(x_{t};\hat{x}_{t|t-1},P_{t|t-1})}-\ell(x_{t},y_{t})\right\}}{\partial x_{t}^{2}}\right|_{x_{t}=\hat{x}_{t|t}^{\mathrm{MAP}}}.$
This method works quite well in most scenarios. Besides this initialization trick, other methods such as Cholesky decomposition [28] or square-root parameterization [29] can also be leveraged to ensure the positive definiteness of the covariance matrix.
An important consideration for the iterative scheme is the stopping criterion. We propose using the KL divergence between two consecutive Gaussian distributions, specifically $\mathcal{N}^{(i)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i)},P_{t|t}^{(i)})$ and $\mathcal{N}^{(i+1)}=\mathcal{N}(x_{t};\hat{x}_{t|t}^{(i+1)},P_{t|t}^{(i+1)})$ to determine when to stop the iteration. As suggested by [14], the stopping condition can be defined as
$D_{\mathrm{KL}}(\mathcal{N}^{(i)}\|\mathcal{N}^{(i+1)})<\gamma,$ (31)
where $\gamma$ is a predefined threshold. This approach is more effective compared to using $D_{\mathrm{KL}}(\mathcal{N}^{(i+1)}\|\mathcal{N}^{(i)})<\gamma$ because $\mathcal{N}^{(i+1)}$ is generally more concentrated. By using $D_{\mathrm{KL}}(\mathcal{N}^{(i)}\|\mathcal{N}^{(i+1)})$, we better capture meaningful changes in the iteration and ensure that the criterion is sensitive enough to detect convergence, while avoiding premature termination of the algorithm.
Recall that in Section III-B, we proved that the prediction step of the moment-matching KF algorithms essentially follows the optimal solution of Gaussian filtering. By combining this step with our natural gradient descent update step (28), we developed a new iterative filter. To emphasize that natural gradient descent is our key contribution, we call it the Natural grAdient GaussiaN apprOximation filter, or NANO filter for short. The pseudocode of the NANO filter is summarized in Algorithm 1. Note that all the expectation computations appearing in Algorithm 1 are suggested to use the efficient unscented transform [30].
Algorithm 1 NANO Filter
Input: Stopping thresohold $\gamma$
Initialization: State estimate $\hat{x}_{0|0}$ and covariance $P_{0|0}$
for each time step $t$ do
     Predict:
     Calculate predicted state mean $\hat{x}_{t|t-1}$ and covariance $P_{t|t-1}$ using (13)
     Update:
     Obtain the noisy measurement $y_{t}$
     Initialize the state estimate $\hat{x}_{t}^{(0)}$ and covariance $P_{t}^{(0)}$ using (30)
     for each iteration number $i$ do
         if (31) is not satisfied then
              Update state estimate and covariance using (28)
         end if
     end for
     $\hat{x}_{t|t}=\hat{x}_{t}^{(i)},P_{t|t}=P_{t}^{(i)}$
end for
V Theoretical Analysis
In this section, we will answer Q3, i.e., we will provide the stability and convergence analysis for NANO filter.
V-A Convergence Analysis
After deriving the update scheme of natural gradient Gaussian filtering in (26), a key question arises: does it converge to the optimal solution of (9b)? The following theorem confirms the local convergence of our update scheme:
Theorem 1.
Consider the Taylor series expansion that is second order in $\delta\hat{x}_{t}$ and first order in $\delta P_{t}^{-1}$. Under this approximation, the iterative update in (26) guarantees convergence, i.e.,
$J_{t}^{(i+1)}\approx$ $\ J_{t}^{(i)}+\left.\frac{\partial{J}}{\partial{\hat{x}_{t}^{\top}}}\right|_{v^{(i)}}\delta\hat{x}_{t}+\frac{1}{2}\left(\delta\hat{x}_{t}\right)^{\top}\left(\left.\frac{\partial^{2}{J}}{\partial{\hat{x}_{t}^{2}}}\right|_{v^{(i)}}\right)\delta\hat{x}_{t}$ (32)
$+\mathrm{Tr}\left(\left.\frac{\partial J}{\partial P_{t}^{-1}}\right|_{v^{(i)}}\delta P_{t}^{-1}\right)$
$\leq$ $\ J_{t}^{(i)},$
where $J_{t}^{(i)}$ is the update cost at the $i$-th iteration, defined as $J_{t}^{(i)}\triangleq J(\hat{x}_{t}^{(i)},P_{t}^{(i)})$. Moreover, equality in (32) holds if and only if $\delta\hat{x}_{t}=0$ and $\delta P_{t}^{-1}=0$.
This theorem indicates that the natural gradient descent iteration in the update step of the NANO filter provides a guarantee of local convergence. This guarantee is achieved by approximating the objective function with second-order accuracy around the mean and first-order accuracy around the inverse of the covariance matrix.
The proof of Theorem 1 is provided in Appendix -B. The key idea is to show that the difference in update costs between consecutive iterations can be expressed as a semi-negative definite quadratic form. Technically, achieving this semi-negative definiteness relies on the use of the Fisher information matrix, which corrects the gradient direction to provide the steepest descent on the Gaussian manifold. This adjustment makes the gradient ”natural” because it aligns with the geometry of the Riemannian space of Gaussian parameters.
For linear Gaussian systems, the NANO filter achieves the optimal solution of equation (9b) in a single iteration, as stated in the following corollary:
Corollary 1.
For linear Gaussian systems in Example 2, the update rule given by (26) converges to the optimal solution of (9b) within one iteration. In other words, a single iteration of NANO filter is equivalent to KF.
The proof of this corollary is provided in Appendix -C. Corollary 1 holds regardless of the initialization of the Gaussian parameters $\hat{x}^{(0)}_{t}$ and $P^{(0)}_{t}$, supporting the fact that the natural gradient is the steepest descent direction in the Gaussian manifold.
V-B Stability Analysis
The stability analysis of our proposed NANO filter is based on the state-space model (1). We focus on the case where the process and measurement noises are both zero-mean Gaussian noise, satisfying $\mathbf{\xi}_{t}\sim\mathcal{N}(0,Q_{t})$ and $\mathbf{\zeta}_{t}\sim\mathcal{\mathcal{missing}}{N}(0,R_{t})$. The measurement-dependent loss $\ell(x_{t},y_{t})$ in the analysis is chosen to be the log-likelihood loss function that satisfies
$\ell(x_{t},y_{t})$ $=-\log p(y_{t}|x_{t})$
$=C_{0}+\frac{1}{2}(y_{t}-g(x_{t}))^{\top}R_{t}^{-1}(y_{t}-g(x_{t})).$
Here, $C_{0}>0$ is a constant value irrelevant to the state and measurement. For the purpose of stability analysis, we make the following regularity assumptions on the system functions $f$ and $g$.
Assumption 2.
The derivatives of the functions $f$ and $g$ are bounded. Specifically, there exists a constant $C>0$ such that
$\left|\frac{\partial f^{i}}{\partial x^{j}}\right|,\left|\frac{\partial g^{k}}{\partial x^{j}}\right|,\left|\frac{\partial^{2}g^{k}}{\partial x^{i}\partial x^{j}}\right|\leq C<\infty,$ (33)
for all $x\in\mathbb{R}^{n}$, $1\leq i,j\leq n$, $1\leq k\leq m$. Here, $x^{i}$ represents the $i$-th component of the state vector $x$; similarly, $f^{i}$ and $g^{k}$ denote the corresponding components of the functions $f$ and $g$, respectively.
Using the standardization of the state $x_{t}$ under the Gaussian distribution, we can rewrite the stationary point condition for update step (18) in the following tensor form:
$\hat{x}_{t|t}^{i}=$ $\hat{x}_{t|t-1}^{i}+(P_{t|t-1})^{ij}\int(R_{t}^{-1})_{kl}\frac{\partial g^{k}}{\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)$ (34a)
×$\times(y_{t}^{l}-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
$(P_{t|t}^{-1})_{ij}=$ $(P_{t|t-1}^{-1})_{ij}$ (34b)
$+\int(R_{t}^{-1})_{kl}\biggl{[}\frac{\partial g^{k}}{\partial x^{i}}(\hat{x}_{t|t}+S_{t|t}z)\frac{\partial g^{l}}{\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)$
$\qquad-\frac{\partial^{2}g^{k}}{\partial x^{i}\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)(y_{t}^{l}-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\biggr{]}$
×$\qquad\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
where $S_{t|t}S_{t|t}^{\top}=P_{t|t}$ is obtained by Cholesky decomposition, and we use the Einstein summation convention [31].
Next, we will substitute the measurement model, $y_{t}^{l}=g^{l}(x_{t})+\zeta_{t}^{l}$ into the tensor form of the stationary point condition in (34), and consider the Taylor expansion of $g(x_{t})$ at $\hat{x}_{t|t}$. For the stationary point condition of $\hat{x}_{t|t}$, let us first define the auxiliary function $h_{j}^{kl}(x)$, for $1\leq j\leq n$ and $1\leq k,l\leq n$:
$h_{j}^{kl}(x)$ $=\int\frac{\partial g^{k}}{\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)$
×$\times(g^{l}(x)-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
then (34a) can be rewritten as
$\hat{x}_{t|t}^{i}=$ $\hat{x}_{t|t-1}^{i}+(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}h_{j}^{kl}(x_{t})$ (35)
$+(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial g^{k}}{\partial x^{j}}(x_{t})\biggr{\}}\zeta_{t}^{l}$
$=$ $\hat{x}_{t|t-1}^{i}+(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}$
×$\times\biggl{[}h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(x_{t}^{q}-\hat{x}_{t|t}^{q})+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr{]}$
$+(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial g^{k}}{\partial x^{j}}(x_{t})\biggr{\}}\zeta_{t}^{l},$
where
[…]
for some $\kappa_{1},\epsilon_{1}>0$, are high-order terms in the Taylor expansion of $h^{kl}_{j}$.
Let us denote the estimation error as $e_{t|t}=x_{t}-\hat{x}_{t|t}$ and $e_{t|t-1}=x_{t}-\hat{x}_{t|t-1}$. Regardless of the specific methods used in the prediction step, the prediction error $e_{t|t-1}$ can be expressed as
$e_{t|t-1}^{i}=$ $f^{i}(x_{t-1})+\xi_{t}^{i}-\hat{x}_{t|t-1}^{i}$ (36)
$=$ $\frac{\partial f^{i}}{\partial x^{j}}(\hat{x}_{t-1|t-1})(x_{t-1}^{j}-\hat{x}_{t-1|t-1}^{j})$
$+\tilde{\psi}^{i}(x_{t-1}-\hat{x}_{t-1|t-1})+\xi_{t}^{i},$
where
$\tilde{\psi}^{i}(x_{t-1}-\hat{x}_{t-1|t-1})$ $\leq\kappa_{2}(\|x_{t-1}-\hat{x}_{t-1|t-1}\|^{2}),$
[…]
for some $\kappa_{2},\epsilon_{2}>0$, which are higher-order terms in the Taylor expansion of $f^{i}$. Subtracting $x_{t}$ from both sides of (35) and applying (36), we obtain
$e_{t|t}^{i}=$ $\frac{\partial f^{i}}{\partial x^{j}}(\hat{x}_{t-1|t-1})e_{t-1|t-1}^{j}+\tilde{\psi}^{i}(x_{t-1}-\hat{x}_{t-1|t-1})$ (37)
$+\xi_{t}^{i}-(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}$
×$\times\biggl{[}h_{j}^{kl}(\hat{x}_{t|t})+\frac{\partial h_{j}^{kl}}{\partial x^{q}}(\hat{x}_{t|t})e_{t|t}^{q}+\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})\biggr{]}$
$-(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial g^{k}}{\partial x^{j}}(x_{t})\biggr{\}}\zeta_{t}^{l}.$
In matrix form, (37) becomes
$e_{t|t}=F_{t-1}e_{t-1|t-1}-H_{t}e_{t|t}-\bar{h}_{t}+\Psi_{t}+\xi_{t}-G_{t}\zeta_{t},$
where $F_{t}$, $H_{t}$, $\bar{h}_{t}$ and $G_{t}$ are matrix- or vector-valued functions with components given by
$(F_{t-1})^{i}_{j}$ $=\frac{\partial f^{i}}{\partial x^{j}}(\hat{x}_{t-1|t-1}),$
$(H_{t})_{q}^{i}$ $=(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\frac{\partial h_{j}^{kl}}{\partial x^{q}}(\hat{x}_{t|t}),$
$\bar{h}_{t}^{i}$ $=(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}h_{j}^{kl}(\hat{x}_{t|t}),$
$(G_{t})_{l}^{i}$ $=(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial g^{k}}{\partial x^{j}}(x_{t})\biggr{\}},$
and
$\Psi_{t}^{i}$ $=\tilde{\psi}^{i}(e_{t-1|t-1})-(P_{t|t-1})^{ij}(R_{t}^{-1})_{kl}\psi_{j}^{kl}(x_{t}-\hat{x}_{t|t})$
are the high-order terms. Therefore, we have
$e_{t|t}=$ $(I+H_{t})^{-1}F_{t-1}e_{t-1|t-1}-(I+H_{t})^{-1}\bar{h}_{t}$ (38)
$+(I+H_{t})^{-1}\Psi_{t}+(I+H_{t})^{-1}\left(\xi_{t}-G_{t}\zeta_{t}\right).$
Similarly, for the stationary point condition of $P_{t|t}$, let us define the auxiliary function $\tilde{h}_{ij}^{kl}(x)$, for $1\leq i,j\leq n$ and $1\leq k,l\leq n$:
$\tilde{h}_{ij}^{kl}(x)=$ $\int\biggl{[}\frac{\partial g^{k}}{\partial x^{i}}(\hat{x}_{t|t}+S_{t|t}z)\frac{\partial g^{l}}{\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)$
$-\frac{\partial^{2}g^{k}}{\partial x^{i}\partial x^{j}}(\hat{x}_{t|t}+S_{t|t}z)(g^{l}(x)-g^{l}(\hat{x}_{t|t}+S_{t|t}z))\biggr{]}$
×$\times\frac{1}{(2\pi)^{\frac{n}{2}}}e^{-\frac{1}{2}\|z\|^{2}}\,\mathrm{d}z,$
then (34b) can be rewritten as
$(P_{t|t}^{-1})_{ij}=$ $(P_{t|t-1}^{-1})_{ij}+(R_{t}^{-1})_{kl}\tilde{h}_{ij}^{kl}(x_{t})$
$-(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial^{2}g^{k}}{\partial x^{i}\partial x^{j}}(x_{t})\biggr{\}}\zeta_{t}^{l}$
$=$ $(P_{t|t-1}^{-1})_{ij}+(R_{t}^{-1})_{kl}\biggl{[}\tilde{h}_{ij}^{kl}(\hat{x}_{t|t})$
$+\frac{\partial\tilde{h}_{ij}^{kl}}{\partial x^{q}}(\hat{x}_{t|t})(x_{t}^{q}-\hat{x}_{t|t}^{q})+\varphi_{ij}^{kl}(x_{t}-\hat{x}_{t|t})\biggr{]}$
$-(R_{t}^{-1})_{kl}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t|t},P_{t|t})}\biggl{\{}\frac{\partial^{2}g^{k}}{\partial x^{i}\partial x^{j}}(x_{t})\biggr{\}}\zeta_{t}^{l},$
where
[…]
for some $\kappa_{3},\epsilon_{3}>0$, are high-order terms in the Taylor expansion of $\tilde{h}^{kl}_{ij}$.
The main idea of stability analysis is based on the application of Lyapunov functions, just as in the case of Kalman-type nonlinear filter [32, 33]. In order to construct the Lyapunov function, let us first introduce the auxiliary covariance matrix $\tilde{P}_{t|t}$ and $\tilde{P}_{t|t-1}$, which evolve according to
$(\tilde{P}_{t|t}^{-1})_{ij}=$ $(\tilde{P}_{t|t-1}^{-1})_{ij}+(R_{t}^{-1})_{kl}\tilde{h}_{ij}^{kl}(\hat{x}_{t|t}),$
$(\tilde{P}_{t|t-1})^{ij}=$ $\frac{\partial f^{i}}{\partial x^{k}}(\hat{x}_{t-1|t-1})\frac{\partial f^{j}}{\partial x^{l}}(\hat{x}_{t-1|t-1})$
×$\times(\tilde{P}_{t-1|t-1})^{kl}+(Q_{t})^{ij},$
and in matrix form
$\tilde{P}_{t|t}^{-1}$ $=\tilde{P}_{t|t-1}^{-1}+D_{t},$ (39)
$\tilde{P}_{t|t-1}$ $=F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t},$
where $D_{t}$ is the matrix with each component $(D_{t})_{ij}=(R_{t}^{-1})_{kl}\tilde{h}_{ij}^{kl}(\hat{x}_{t|t})$. Note that the evolution of $\tilde{P}_{t|t}$ and $\tilde{P}_{t|t-1}$ does not depend directly on the errors $e_{t|t}$, $e_{t|t-1}$ or the noise terms $\xi_{t}$, $\zeta_{t}$. This makes the following positive definiteness and boundedness assumption largely a condition on the system itself, much like the detectable and controllable conditions for linear systems [34].
Assumption 3.
There exist constants $\underline{p},\bar{p}>0$, such that
$0<\underline{p}I\leq\tilde{P}_{t|t}\leq\bar{p}I<\infty,\ \forall\ t\geq 0,$
$0<\underline{p}I\leq\tilde{P}_{t|t-1}\leq\bar{p}I<\infty,\ \forall\ t\geq 0$
Our main stability result is stated in the following theorem. Generally speaking, this theorem proves the stability of our proposed method for those systems with almost linear measurement functions and small noise.
Theorem 2.
Under Assumption 2 and 3, the estimation error $e_{t|t}$ is exponentially bounded in the mean square for systems with almost linear measurement functions, i.e., there exist $\epsilon,\epsilon^{\prime},\lambda>0$, such that
$\mathbb{E}\left\|e_{t|t}\right\|^{2}\leq\epsilon\|e_{0|0}\|^{2}\biggl{(}\frac{1}{1+\lambda}\biggr{)}^{t}+\epsilon^{\prime},\ \forall\ t\geq 0,$
as long as the initial error and the strength of the system noise are small enough, that is $\|e_{0|0}\|\leq\delta$, $\mathbb{E}\left\{\xi_{t}\xi_{t}^{\top}\right\}\leq\delta I$ and $\mathbb{E}\left\{\zeta_{t}\zeta_{t}^{\top}\right\}\leq\delta I$ for some given $\delta>0$.
Proof.
Utilizing (38) and (39), $e_{t|t}^{\top}\tilde{P}_{t|t}^{-1}e_{t|t}$ can be computed as follows:
$e_{t|t}^{\top}\tilde{P}_{t|t}^{-1}e_{t|t}$ (40)
$=$ $e_{t-1|t-1}^{\top}F_{t-1}^{\top}(I+H_{t})^{-1}$
×$\quad\quad\times\left[\left(F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t}\right)^{-1}+D_{t}\right]$
×$\quad\quad\times(I+H_{t})^{-1}F_{t-1}e_{t-1|t-1}$
$\quad-2e_{t-1|t-1}^{\top}F_{t-1}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\bar{h}_{t}$
$\quad+\bar{h}_{t}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\bar{h}_{t}$
$\quad+\Psi_{t}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\Psi_{t}$
$\quad+\left(\xi_{t}-G_{t}\zeta_{t}\right)^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\xi_{t}-G_{t}\zeta_{t}\right)$
$\quad+2e_{t-1|t-1}F_{t-1}^{\top}(I+H_{t})^{-1}$
×$\qquad\times\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\Psi_{t}+\xi_{t}-G_{t}\zeta_{t}\right)$
$\quad-2\bar{h}_{t}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\Psi_{t}+\xi_{t}-G_{t}\zeta_{t}\right)$
$\quad+2\Psi_{t}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\left(\xi_{t}-G_{t}\zeta_{t}\right).$
For the first term on the right-hand side of (40), with the fact that the measurement function $h(x)$ is almost linear, there exists $\lambda_{1}>0$ such that
$F_{t-1}^{\top}$ $(I+H_{t})^{-1}\left[\left(F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t}\right)^{-1}+D_{t}\right]$ (41)
×$\times(I+H_{t})^{-1}F_{t-1}\leq\frac{1}{1+\lambda_{1}}\tilde{P}_{t-1|t-1}.$
In fact, if $g(x)=Cx$ is a linear function with some constant matrix $C\in\mathbb{R}^{m\times n}$, then
$H_{t}=P_{t|t-1}C^{\top}R_{t}^{-1}C,$
$D_{t}=C^{\top}R_{t}^{-1}C,$
and $\tilde{P}_{t|t}=P_{t|t}$, $\tilde{P}_{t|t-1}=P_{t|t-1}$. In this way, we have
$F_{t-1}^{\top}(I+H_{t})^{-1}\left[\left(F_{t-1}\tilde{P}_{t-1|t-1}F_{t-1}^{\top}+Q_{t}\right)^{-1}+D_{t}\right]$
×$\qquad\times(I+H_{t})^{-1}F_{t-1}$
$=$ $F_{t-1}^{\top}(I+P_{t|t-1}C^{\top}R_{t}^{-1}C)^{-1}P_{t|t-1}^{-1}F_{t-1}$
$=$ $F_{t-1}^{\top}(I+P_{t|t-1}C^{\top}R_{t}^{-1}C)^{-1}(F_{t-1}^{\top}P_{t-1|t-1}F_{t-1}^{\top})^{-1}$
×$\quad\times(I+(F_{t-1}P_{t-1|t-1}F_{t-1}^{\top})^{-1}Q_{t})^{-1}F_{t-1}.$
By observing that
$(I+P_{t|t-1}C^{\top}R_{t}^{-1}C)^{-1}$ $\leq\frac{1}{1+\lambda}I,$
$(I+(F_{t-1}P_{t-1|t-1}F_{t-1}^{\top})^{-1}Q_{t})^{-1}$ $\leq\frac{1}{1+\lambda^{\prime}}I,$
with $\lambda=\lambda_{\min}(P_{t|t-1}C^{\top}R_{t}^{-1}C)>0$ and $\lambda^{\prime}=\lambda_{\min}((F_{t-1}P_{t-1|t-1}F_{t-1}^{\top})^{-1}Q_{t})>0$, therefore (41) holds for some $\lambda_{1}>0$ for linear systems. Because of the continuity of (41) with respect to the observation function $g$, (41) also holds for those systems with almost linear measurement functions.
Other terms on the right-hand side of (40) can also be bounded. Firstly, there exists $\delta_{1}>0$, such that
$\bar{h}_{t}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\bar{h}_{t}\leq\delta_{1}.$ (42)
Because $\bar{h}_{t}\equiv 0$ for linear observation functions, the constant $\delta_{1}$ above can be chosen to be sufficiently small for almost linear observation functions.
Secondly, there exist $\delta_{2}>0$, such that
$-2e_{t-1|t-1}^{\top}F_{t-1}^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}\bar{h}_{t}\leq\delta_{2}\|e_{t-1|t-1}\|.$ (43)
Finally, because $\xi_{t}$, $\zeta_{t}$ are independent Gaussian random variables, there exists $\delta_{3}>0$, such that
$\mathbb{E}\left\{(\xi_{t}-G_{t}\zeta_{t})^{\top}(I+H_{t})^{-1}\tilde{P}_{t|t}^{-1}(I+H_{t})^{-1}(\xi_{t}-G_{t}\zeta_{t})\right\}\leq\delta_{3}.$ (44)
Other terms on the right-hand side of (40) are related to the high order terms $\Psi_{t}$. Therefore, combining (41), (42) to (44), there exists $\eta>0$, such that if $\|e_{t|t}\|<\eta$ for all $t\geq 0$, then
$\mathbb{E}\left\{\left.e_{t|t}^{\top}\tilde{P}_{t|t}^{-1}e_{t|t}\right|e_{t-1|t-1}\right\}$
$\leq$ $\frac{1}{1+\lambda_{0}}e_{t-1|t-1}^{\top}\tilde{P}_{t-1|t-1}^{-1}e_{t-1|t-1}+\delta_{2}\cdot\mathbb{E}\|e_{t-1|t-1}\|+\delta,$
for some $0<\lambda_{0}<\lambda_{1}$.
With the fact that $\underline{p}I\leq\tilde{P}_{t|t}\leq\bar{p}I$, there exist $\epsilon_{1},\epsilon_{2}>0$, such that
$\mathbb{E}\left\{\left.(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right|e_{t-1|t-1}\right\}$
$\leq$ $\frac{1}{1+\lambda_{0}}(\|e_{t-1|t-1}\|-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(\|e_{t-1|t-1}\|-\epsilon_{1})+\epsilon_{2},$
as long as $\|e_{t-1|t-1}\|<\eta$.
Also, if there exists $\tilde{\eta}>0$, such that $\tilde{\eta}<\|e_{t-1|t-1}\|<\eta$, for sufficiently small $\epsilon_{1},\epsilon_{2}$, we have the following supermartingale-like property:
$\mathbb{E}\left\{\left.(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right|e_{t-1|t-1}\right\}$ (45)
$-(\|e_{t-1|t-1}\|-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(\|e_{t-1|t-1}\|-\epsilon_{1})$
$\leq$ $-\frac{\lambda_{0}}{2(1+\lambda_{0})\bar{p}}\tilde{\eta}^{2}+\epsilon_{2}<0.$
Hence, if the initial estimation error $\|e_{0|0}\|<\eta$, then we can recursively use (V-B) to compute the estimation error at time $t$,
$\mathbb{E}\left\{(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right\}$
$\leq$ $\epsilon_{2}\sum_{k=0}^{t-1}\frac{1}{(1+\lambda_{0})^{k}}+\frac{1}{\underline{p}}\|e_{0|0}-\epsilon_{1}\|^{2}\left(\frac{1}{1+\lambda_{0}}\right)^{t}$
$\leq$ $\frac{1+\lambda_{0}}{\lambda_{0}}\epsilon_{2}+\frac{1}{\underline{p}}\|e_{0|0}-\epsilon_{1}\|^{2}\left(\frac{1}{1+\lambda_{0}}\right)^{t}.$
Since
$\frac{1}{\bar{p}}\|\|e_{t|t}\|-\epsilon_{1}\|^{2}\leq(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1}),$
and thus,
$\|e_{t|t}\|^{2}\leq\left(\sqrt{\bar{p}(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})}+\epsilon_{1}\right)^{2},$
for sufficiently small $\epsilon_{1}$, we obtained the exponentially bounded in mean square for $e_{t|t}$, i.e., there exist $\epsilon,\epsilon^{\prime}>0$, such that
$\mathbb{E}\|e_{t|t}\|^{2}\leq\epsilon\|e_{0|0}\|^{2}\biggl{(}\frac{1}{1+\lambda_{0}}\biggr{)}^{t}+\epsilon^{\prime},\ \forall\ t\geq 0.$
VI Robustifying NANO Filter with Gibbs Posterior
As illustrated in (6), the objective of update step of Bayesian filtering is to balance the information provided by the measurement data with that of the prior distribution. Nevertheless, acquiring the information from the measurement data necessitates a complete probabilistic model of the measurement-generating process. In other words, a perfect model of the output distribution $p(y_{t}|x_{t})$ must be known in advance. Unfortunately, specifying the true model of $p(y_{t}|x_{t})$ is often impossible due to sensor malfunctions or overlooked dynamics [35, 36], which introduces outliers in the measurement data. In this case, a viable solution is to replace the likelihood function term in (6) to a general measurement-dependent risk function, which leads to the following optimization problem:
$p_{\mathrm{G}}(x_{t}|y_{1:t})=$ $\operatorname*{arg\,min}_{q(x_{t})}\Big{\{}\mathbb{E}_{q(x_{t})}\left\{\ell^{G}(x_{t},y_{t})\right\}$ (46)
$+D_{\mathrm{KL}}\left\{q(x_{t})||p(x_{t}|y_{1:t-1})\right\}\Big{\}}.$
Here, $\ell^{G}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is the generalized measurement-dependent loss function. It turns out that the solution to the optimization problem (46) is known as the Gibbs posterior, which has an analytical form akin to the standard Bayesian posterior:
$p_{\mathrm{G}}(x_{t}|y_{1:t})=\frac{\exp\{\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})}{\int\exp\{\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}}.$
Here, we denote the Gibbs posterior as $p_{\mathrm{G}}(x_{t}|y_{1:t})$ to differentiate from the standard Bayesian posterior $p(x_{t}|y_{1:t})$. In fact, the standard Bayesian posterior can be regarded as a special case of the Gibbs posterior by setting the loss function as the negation of the logarithm of the likelihood function, i.e., $\ell^{G}(x_{t},y_{t})=-\log p(y_{t}|x_{t})$.
Remark 2.
A notable observation is that the analysis of the Gaussian approximation in Section III-C and the derivation of the NANO filter in Section IV do not depend on the specific form of the loss function $\ell(x_{t},y_{t})$. Therefore, since Gibbs Bayesian filtering in (46) only modifies the loss function $\ell(x_{t},y_{t})$ in (6) to a more general form, $\ell^{G}(x_{t},y_{t})$, all of the previous analysis and algorithmic design can naturally extend to this case.
Then, we provide several potential choices of the measurement-dependent loss functions that can tackle model misspecifications:
Choice 1: Composite Likelihood: One popular choice is to use the composite likelihood loss, i.e., combine multiple likelihood functions to achieve robust and adaptive performance under varying conditions. For example, for single-dimension measurement case, we can assume that the loss function is a composition of Gaussian likelihood and Laplace likelihood:
$\ell^{\mathrm{h}}(x_{t},y_{t})$
$=$ $\begin{cases}-\log{\mathcal{N}(y_{t};h(x_{t}),1)}&\text{if}\quad|y_{t}-h(x_{t})|\leq\delta,\\ -\log{\mathrm{Laplace}(y_{t};h(x_{t}),\frac{1}{\delta})}&\text{otherwise.}\end{cases}$
=$\stackrel{{\scriptstyle c}}{{=}}$$\begin{cases}\frac{1}{2}|y_{t}-h(x_{t})|^{2}&\text{if}\quad|y_{t}-h(x_{t})|\leq\delta,\\ \delta\left(|y_{t}-h(x_{t})|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$
Here, the notation ``=$\stackrel{{\scriptstyle c}}{{=}}"$ represents that two expressions are equivalent up to an additive constant, and $\delta>0$ is the threshold variable. In fact, $\ell^{\mathrm{h}}(x_{t},y_{t})$ represents the well-known Huber loss [37] used in robust regression, which is known to be less sensitive to outliers in data than the squared error loss. In practical applications, we utilize a smooth approximation of the Huber loss, known as the Pseudo-Huber loss, to improve optimization efficiency and ensure differentiability. The Pseudo-Huber loss is defined as:
$\ell^{\mathrm{ph}}(x_{t},y_{t})=\delta^{2}\left(\sqrt{1+\left(y_{t}-h(x_{t})\right)^{2}/{\delta}^{2}}-1\right).$ (47)
Choice 2: Weighted Log-Likelihood: Besides the composition of different likelihoods, one natural choice to achieve a robust loss function is to scale the log-likelihood loss function with a data-dependent weighting term:
$\ell^{\mathrm{w}}(x_{t},y_{t})=-w(x_{t},y_{t})\cdot\log{p(y_{t}|x_{t})}.$ (48)
Here, $w:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}_{+}$ is the weighting function. The philosophy behind this form of loss is that reweighting the effect of outliers in the filtering procedure can potentionally improve robustness. The choices of the weighting function can be inspired from several domains. For example, we can use the so-called inverse multi-quadratic weighting function [38]
$w(x_{t},y_{t})=(1+\left\|y_{t}-h(x_{t})\right\|_{R_{t}^{-1}}^{2}/c^{2})^{-1},$
where $c>0$ is a constant number. Other choices of the weighting function can also be found in Section 3.3 of [39].
Remark 3.
When $w(x_{t},y_{t})$ is chosen as the constant value that is smaller than 1, it can be regarded as the result of performing the so-called exponential density rescaling for convlutional Bayes filter [40], which already shows robustness for systems with outliers. In this case, $w(x_{t},y_{t})$ can be regarded as the Lagrange multiplier which balances the compression and recontruction in the well-known information bottleneck problem [40].
Choice 3: Divergence-Dependent Loss Function: The final option is to utilize robust divergence to fit the data. This approach is based on the fact that minimizing the negative log-likelihood essentially amounts to minimizing the KL divergence between the true data distribution $p_{\mathrm{true}}(y_{t})$, which includes outliers, and the assumed likelihood. More specifically, the minimization of the measurement-dependent loss in (46) can be derived as follows:
$\operatorname*{arg\,min}_{q(x_{t})}\mathbb{E}_{q(x_{t})}\left\{D_{\mathrm{KL}}\left(p_{\mathrm{true}}(y_{t})\|p(y_{t}|x_{t})\right)\right\}$ (49)
$=$ $\operatorname*{arg\,min}_{q(x_{t})}\mathbb{E}_{q(x_{t})}\mathbb{E}_{p_{\mathrm{true}}(y_{t})}\left\{-\log{p(y_{t}|x_{t})}\right\}$
$\approx$ $\operatorname*{arg\,min}_{q(x_{t})}\mathbb{E}_{q(x_{t})}\left\{-\log{p(y_{t}|x_{t})}\right\}.$
In this equation, the first equality arises because the term $p_{\mathrm{true}}(y_{t})$ is treated as a constant regarding $q(x_{t})$ and thus omitted, since the objective is to find the minimizer, not the minimum value. The final approximation holds because we cannot directly access the true data distribution; instead, we use the sample $y_{t}$ to approximate the expected value.
Based on this understanding, one natural idea for designing a robust loss is to replace the KL divergence with a more robust divergence, such as $\beta$-divergence or $\gamma$-divergence. For instance, replacing the KL divergence with $\beta$-divergence in (49) leads to the $\beta$ loss function $\ell^{\beta}(x_{t},y_{t})$ [35]:
$\ell^{\beta}(x_{t},y_{t})=-\frac{\beta+1}{\beta}p(y_{t}|x_{t})^{\beta}+\int{p(y|x_{t})^{\beta+1}\mathrm{d}y}.$ (50)
Remark 4.
Formally analyzing which loss function provides superior robustness is challenging, as each is grounded in different principles. A common approach is to apply Huber’s robust statistics theory [41], which assesses the robustness of an estimator using its influence function. An estimator is considered robust if its influence function remains bounded as an outlier’s value increases indefinitely. Previous studies have conducted case-specific analyses of robustness for Huber loss [42], weighted loss [39], and $\beta$ loss [35] in linear Gaussian systems, demonstrating that adjusting the loss function can improve robustness. Since this paper focuses on applying robust loss functions to enhance the NANO filter, a detailed formal analysis is left for future work.
VII Discussions
Natural Gradient Descent for Gaussian Approximation: The use of natural gradient descent for finding optimal Gaussian approximations is well-documented in the literature, with early works dating back to [43] with applications for Gaussian process regression. After that, related methods have been explored in various domains, including robot batch estimation [44, 45], Bayesian deep learning [46, 47], approximate inference [48], optimization [49] and more.
We utilize the natural gradient method to find the optimal Gaussian approximation for Bayesian filtering in this paper. Theoretically, the advantage of incorporating the natural gradient in Gaussian filtering may arise from the profound geometric properties of the statistical manifolds generated by the family of Gaussian distributions, as discussed in [50, 51]. Intuitively, the computation of the natural gradient involves second-order derivatives of the probability density functions. Thus, geometric concepts such as Riemannian metrics, curvatures, and geodesics on the statistical manifold can be engaged and reflected in the algorithm, leading to better numerical performance. Nevertheless, a precise description of the relationship between the natural gradient and Gaussian distribution remains an important research direction for further exploration.
Gradient-Based Gaussian Filters: Gradient-based Gaussian filters, which use gradient descent and its variants to solve Gaussian filtering, were initially proposed in [52], utilizing Monte Carlo techniques to construct gradient descent information. Subsequently, cubature rules were introduced to replace the Monte Carlo method, and various optimization approaches, such as the conditional gradient method and the alternating direction method of multipliers, were employed to replace basic gradient descent [53, 54, 55].
Unfortunately, these methods have certain assumptions and limitations compared to our approach. For instance, (i) the methods in [52, 53, 54] are restricted to systems with Gaussian noise; (ii) [52] and [55] require performing linear approximations of the measurement model; and (iii) [53] modifies the implicit hard constraints of the original problem into soft constraints, rendering it inequivalent to the original problem. Also, they both lack rigorous proofs of algorithm convergence and stability guarantees for the estimation error, which we provide in Section V. Additionally, their optimization problems are constructed based on direct minimization of forward KL divergence, which does not support the extension to the Gibbs posterior as proposed in our method.
Also, we would like to emphasize that, although these works employ gradient methods for Gaussian approximation, they do not theoretically analyze the stationary points corresponding to the optimal Gaussian approximation. For example, they do not discuss the rationale for using gradient-based methods in the update process, nor do they consider the Gaussian approximation in the prediction step. In contrast, our paper systematically discusses both the prediction and update steps, providing a comprehensive analysis of the stationary points for the optimal Gaussian approximation.
Kalman Filtering as Natural Gradient Descent: Corollary 1 shows that the canonical Kalman Filter (KF) can be interpreted as a single iteration of the natural gradient in our proposed NANO filter. This result is similar to the findings in [56], where the equivalence between the KF and online natural gradient descent is established. However, [56] focuses on parameter estimation, where the state is treated as a deterministic variable representing the parameters of the measurement model. In that context, the natural gradient is applied to maximize the likelihood with respect to the measurement model. In contrast, our work considers the more general linear Gaussian filtering problem, where the state is a random variable, and the natural gradient is used to update the cost with respect to the posterior distribution of the state. Therefore, our result is a generalization of the result in [56].
VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment. All the simulations and experiment are evaluated by root mean square error which is defined by
$\mathrm{RMSE}=\sqrt{\frac{\sum_{t=1}^{T}\|x_{t}-\hat{x}_{t}\|^{2}}{n\cdot T}},$
where $T$ is the total time step for each trajectory. As it is generally impossible to compare with all existing Gaussian filters, we focus on the most popular ones as baselines for our analysis, including the KF, EKF, UKF, IEKF, and PLF. These filters are all mentioned repeatedly in the well-regarded textbook [10].
VIII-A Linear System Simulation: Wiener Velocity Model
First, we perform simulations for the canonical linear Gaussian system, specifically using the Wiener velocity model that is commonly employed for target tracking simulations [10]. In this model, the state represents the position and velocity of a moving object in two dimensions. The state vector is defined as $x=\left[p_{x}\ p_{y}\ v_{x}\ v_{y}\right]^{\top}$, where $p_{x}$ and $p_{y}$ are the object’s positions along the longitude and lateral directions, and $v_{x}$ and $v_{y}$ are the corresponding velocities. The measurements are direct, noisy observations of the position components. The state transition model with a time step $\Delta t=0.1$ and the measurement model can be described as
$x_{t+1}$ $=\begin{bmatrix}1&0&\Delta t&0\\ 0&1&0&\Delta t\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}x_{t}+\xi_{t},$
$y_{t}$ $=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\end{bmatrix}x_{t}+\zeta_{t}.$
Here, $\xi_{t}\sim\mathcal{N}(\xi_{t};0,Q)$ and $\zeta_{t}\sim\mathcal{N}(\zeta_{t};0,R)$ are the process and measurement noises with covariance matrices given by
$Q=\begin{bmatrix}\frac{\Delta t^{3}}{3}&0&\frac{\Delta t^{2}}{2}&0\\ 0&\frac{\Delta t^{3}}{3}&0&\frac{\Delta t^{2}}{2}\\ \frac{\Delta t^{2}}{2}&0&\Delta t&0\\ 0&\frac{\Delta t^{2}}{2}&0&\Delta t\end{bmatrix},R=\mathbb{I}_{2\times 2}.$
Figure 1: Box plot of RMSE of KF, UKF, PLF and NANO filter, for the standard Wiener velocity model.
Figure 2: Box plot of RMSE for KF, UKF, PLF, NANO filter, and its robust variants: $\beta$-NANO filter, Huber-NANO filter, and Weight-NANO filter with various parameter values, for the Wiener velocity model with measurement outliers. Note that the small black square “
$\blacksquare$
” represents the average RMSE over all the Monte Carlo experiments.
To validate our method, we consider two scenarios. In the first scenario, the proposed NANO filter is compared against the KF, the optimal filter for linear Gaussian systems, as well as UKF and PLF. As shown in Fig. 1, the NANO filter achieves the same performance as the KF family within one iteration, supporting the result in Corollary 1 that KF equals a single iteration of NANO filter. In the second scenario, we consider the case where the measurement data deviates from the measurement model due to contamination of outliers. Specifically, the measurement data has a 0.1 probability of being contaminated, modeled as $\zeta_{t}\sim 0.9\cdot\mathcal{N}(0,R)+0.1\cdot\mathcal{N}(0,1000R)$. For this case, as illustrated in Fig. 2, we evaluate the robust variants of the NANO filter, including the Huber-NANO filter, Weight-NANO filter, and $\beta$-NANO filter, which utilize the robust loss functions $\ell^{\mathrm{ph}}$ (47), $\ell^{\mathrm{w}}$ (48), and $\ell^{\beta}$ (50), respectively. The box plots indicate that the robust variants of the NANO filter generally outperform the standard NANO filter using different parameters. Notably, the Weight-NANO filter with $c=25$ shows the best performance, significantly improving estimation accuracy.
VIII-B Nonlinear System Simulation: Air-traffic Control
Next, we perform simulations for a nonlinear system, focusing on an air-traffic control scenario where an aircraft executes a maneuvering turn in a horizontal plane at a height $h=50$ with respect to the radar [14]. The state vector is defined as $x=\left[p_{x},\ v_{x},\ p_{y},\ v_{y},\ \omega\right]^{\top}$, where $p_{x},p_{y}$ are the positions, $v_{x},v_{y}$ are the velocities in the planar coordinates, and $\omega$ is the turn rate. The kinematics of the turning motion, represented by the state transition model, and the measurement model are given by:
$x_{t+1}$ $=\begin{bmatrix}1&\frac{\sin\omega\Delta t}{\omega}&0&-\frac{(1-\cos\omega\Delta t)}{\Omega}&0\\ 0&\cos\omega\Delta t&0&-\sin\omega\Delta t&0\\ 0&\frac{(1-\cos\omega\Delta t)}{\omega}&1&\frac{\sin\omega\Delta t}{\omega}&0\\ 0&\sin\omega\Delta t&0&\cos\omega\Delta t&0\\ 0&0&0&0&1\end{bmatrix}x_{t}+\xi_{t},$
$y_{t}$ $=\begin{bmatrix}\sqrt{p_{x}^{2}+p_{y}^{2}+h^{2}}\\ \operatorname{atan}\left(\frac{p_{y}}{p_{x}}\right)\\ \operatorname{atan}\left(\frac{h}{\sqrt{p_{x}^{2}+p_{y}^{2}}}\right)\\ \frac{p_{x}\dot{p}_{x}+p_{y}\dot{p}_{y}}{\sqrt{p_{x}^{2}+p_{y}^{2}+h^{2}}}\end{bmatrix}+\zeta_{t}.$
Here, $\Delta t=0.2$ is the sampling period. The process noise $\xi_{t}\sim\mathcal{N}(0,Q)$ and measurement noise $\zeta_{t}\sim\mathcal{N}(0,R)$ have covariances given by:
$Q$ $=\begin{bmatrix}q_{1}\cdot\frac{\Delta t^{3}}{3}&q_{1}\cdot\frac{\Delta t^{2}}{2}&0&0&0&\\ q_{1}\cdot\frac{\Delta t^{2}}{2}&q_{1}\cdot\Delta t&0&0&0\\ 0&0&q_{1}\cdot\frac{\Delta t^{3}}{3}&q_{1}\cdot\frac{\Delta t^{2}}{2}&0\\ 0&0&q_{1}\cdot\frac{\Delta t^{2}}{2}&q_{1}\cdot\Delta t&0\\ 0&0&0&0&q_{2}\cdot\Delta t\end{bmatrix},$
$R$ $=\mathrm{diag}\left(\left[\sigma_{r}^{2},\ \sigma_{\varphi}^{2},\ \sigma_{\theta}^{2},\ \sigma_{\dot{r}}^{2}\right]\right).$
Figure 3: Box plot of RMSE for EKF, UKF, IEKF, PLF and NANO filter, for the standard air-traffic control model.
Figure 4: Box plot of RMSE for EKF, UKF, IEKF, PLF, NANO filter, and its robust variants: $\beta$-NANO filter, Huber-NANO filter, and Weight-NANO filter with various parameter values, for the air-traffic model with measurement outliers.
Similar to the linear case, we validate our method under two scenarios for the nonlinear system. In the first scenario, shown in Fig. 3, the NANO filter is compared with the EKF, UKF, IEKF, and PLF. The results demonstrate that the NANO filter significantly outperforms these established filters, achieving nearly a half reduction in RMSE. This substantial improvement highlights the NANO filter’s superior effectiveness in handling nonlinear dynamics. In the second scenario, we introduce measurement outliers to reflect conditions where the measurement data deviates from the assumed model. Specifically, we assume the actual measurement noise follows a mixture model: $\zeta_{t}\sim 0.9\cdot\mathcal{N}(\zeta_{t};0,R)+0.1\cdot\mathcal{N}(\zeta_{t};0,100\cdot R)$, indicating a 10% probability that measurements are contaminated by noise with a tenfold increase in standard deviation. As shown in Fig. 4, robust variants of the NANO filter—including the $\beta$-NANO, Huber-NANO, and Weight-NANO filters—significantly improve estimation accuracy in the presence of measurement outliers. Notably, the $\beta$-NANO filter with $\beta=10^{-2}$ exhibits the best performance, effectively mitigating the impact of outliers.
VIII-C Real-world Experiment: Unmanned Ground Vehicle Localization
We conducted a real-world experiment to demonstrate the effectiveness of our method using an unmanned ground vehicle (UGV) equipped with a Lidar sensor for environmental perception, as shown in Fig. 5. The vehicle’s state is represented by $x=\left[\begin{matrix}p_{x}&p_{y}&\theta\end{matrix}\right]^{\top}$, which includes its 2D position ($p_{x}$ and $p_{y}$) and orientation angle ($\theta$). The longitudinal velocity $v$ and yaw rate $\omega$ serve as control inputs [57]. The Lidar has a 240-degree detection range with 0.33-degree resolution. The UGV was manually controlled throughout the experiment.
Figure 5: (a) The UGV. (b) Experiment field. The three red traffic cones serve as landmarks for positioning.
The vehicle kinematics are modeled as:
$\begin{bmatrix}p_{x,t+1}\\ p_{y,t+1}\\ \theta_{t+1}\end{bmatrix}=\begin{bmatrix}p_{x,t}\\ p_{y,t}\\ \theta_{t}\end{bmatrix}+\begin{bmatrix}v_{t}\cdot\cos\theta_{t}\\ v_{t}\cdot\sin\theta_{t}\\ \omega_{t}\end{bmatrix}\cdot\Delta t+\xi_{t},$
where $\Delta t=0.0667s$ is the sample period, $v_{t}$ is the longitudinal velocity, $\omega_{t}$ is the yaw rate, and $\xi_{t}$ is the process noise. The observation model is given by:
$y_{t}=[d^{1}_{t},d^{2}_{t},d^{3}_{t},\alpha^{1}_{t},\alpha^{2}_{t},\alpha^{3}_{t}]^{\top}+\zeta_{t},$
where $d^{i}$ and $\alpha^{i}$ ($i=1,2,3$) denote the relative distance and orientation angle between the UGV and each traffic cone:
$d^{i}$ $=\|(p^{tc,i}_{x}-p_{x}-l\cos\theta,p^{tc,i}_{y}-p_{y}-l\sin\theta)\|_{2},$ (51)
$\alpha^{i}$ $=\arctan\left(\frac{p^{tc,i}_{y}-p_{y}-l\sin\theta}{p^{tc,i}_{x}-p_{x}-l\cos\theta}\right)-\theta.$
Here $(p^{tc,i}_{x},p^{tc,i}_{y})$ is the position of the $i$-th traffic cone, and $l$ represents the longitudinal installation offset of the Lidar with respect to the robot center, as depicted in Fig. 6.
Figure 6: (a) The diagram of the vehicle’s states containing the 2D position of the vehicle $p_{x}$ and $p_{y}$, and the orientation angle $\theta$. (b) The diagram of the measurement model. Note that the red circle represents the red traffic cone.
Figure 7: Box plot of RMSE for KF, UKF, PLF, NANO and $\beta$-NANO, Huber-NANO, Weight-NANO with different values of the corresponding parameters.
Figure 8: A sample of the real trajectory and its estimates of the UGV estimation.
The experiment began by collecting raw data, including ground truth states, control inputs, and Lidar point clouds. The Lidar data were processed to extract relative distances and angles to the traffic cones, resulting in a dataset comprising trajectories totaling 11 minutes. Using ground truth data obtained from a high-precision motion capture system and the measurement data, we estimated noise parameters, assuming Gaussian distributions with independent noise components. The dataset was then segmented into 50 sub-trajectories, each lasting 6.67 seconds (100 time steps). Figures 7 and 8 compare the RMSE performance of the EKF, UKF, IEKF, PLF, and NANO filters, including their robust variants. The results indicate that the NANO filter consistently outperforms traditional Gaussian filters, achieving higher accuracy. Furthermore, the robust NANO variants demonstrate improved performance, highlighting their effectiveness in handling measurement outliers typically encountered in real-world conditions.
TABLE II: Run Time per time step of different algorithms on our experiments (ms)
Algorithms Wiener Velocity Wiener Velocity (outlier) Air-traffic Control Air-traffic Control (outlier) UGV Localization
EKF 0.023 0.028 1.456 1.439 0.809
UKF 0.105 0.113 0.712 0.659 0.447
iEKF 0.023 0.039 2.028 2.015 2.337
PLF 0.264 0.275 1.023 1.313 0.870
NANO filter 0.406 0.441 1.980 2.023 2.458
Previous results for simulations and experiment demonstrate that the NANO filter significantly improves RMSE performance compared to traditional Gaussian filters. However, a critical aspect to consider is the computational cost per time step for each algorithm. Table II shows that for nonlinear systems, the average runtime of NANO filter is approximately twice that of the EKF, which indicates that despite being an iterative method, NANO maintains high efficiency and quickly converges. This balance of improved accuracy with reasonable computational demands highlights the practicality of the NANO filter, even in real-time applications.
IX Conclusion
In this paper, we address the estimation errors commonly introduced by linearization techniques in Gaussian filters like the EKF and UKF for nonlinear systems. We reformulate the prediction and update steps of Gaussian filtering as optimization problems. While the prediction step mirrors moment-matching filters by calculating the first two moments of the prior distribution, the update step poses more challenges due to its highly nonlinear nature. To overcome these issues, we propose an iterative approach called the NANO filter, which avoids linearization by directly optimizing the update step using the natural gradient derived from the Fisher information matrix. This allows us to account for the curvature of the parameter space and ensures more accurate updates. We prove that the NANO filter converges locally to the optimal Gaussian approximation at each time step and show that it provides exponential error bounds for nearly linear systems with low noise. Experimental results demonstrate that the NANO filter outperforms popular filters like EKF, UKF, and others, while maintaining similar computational efficiency.
-A Proof of Lemma 1
Proof.
The log-likelihood function of a Gaussian distribution $\mathcal{N}(x;\mu,\Sigma)$ is given by
$\log\mathcal{N}(x;\mu,\Sigma)=$ $-\frac{1}{2}\big{[}\log|\Sigma|$
$+(x-\mu)^{\top}\Sigma^{-1}(x-\mu)+\log(2\pi)^{n}\big{]},$
where $n$ is the dimensionality of $x$. Taking the expectation with respect to the probability density function $p(x)$, we obtain
$\mathbb{E}_{p(x)}\left\{\log\mathcal{N}(x;\mu,\Sigma)\right\}=$ $-\frac{1}{2}\big{[}\log|(2\pi)^{n}\Sigma|$
$+\mathbb{E}_{p(x)}\left\{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right\}\big{]}.$
We first compute the gradient of this expectation with respect to $\mu$. The relevant term involving $\mu$ is the quadratic form $(x-\mu)^{\top}\Sigma^{-1}(x-\mu)$. Differentiating this term with respect to $\mu$ gives
$\frac{\partial}{\partial\mu}\mathbb{E}_{p(x)}\left\{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right\}=-2\Sigma^{-1}\mathbb{E}_{p(x)}\left\{x-\mu\right\}.$
Setting this derivative equal to zero, we obtain (11a). Next, we compute the gradient with respect to $\Sigma$. The relevant terms involving $\Sigma$ are the log-determinant $\log|\Sigma|$ and the quadratic form $(x-\mu)^{\top}\Sigma^{-1}(x-\mu)$. The gradient of the log-determinant with respect to $\Sigma$ is $\frac{\partial}{\partial\Sigma}\log|\Sigma|=\Sigma^{-1}$. The gradient of the quadratic form is
$\frac{\partial}{\partial\Sigma}\mathbb{E}_{p(x)}\left\{(x-\mu)^{\top}\Sigma^{-1}(x-\mu)\right\}$
$=$ $-\Sigma^{-1}\mathbb{E}_{p(x)}\left\{(x-\mu)(x-\mu)^{\top}\right\}\Sigma^{-1}.$
Combining these results, we have
$\frac{\partial}{\partial\Sigma}\mathbb{E}_{p(x)}\left\{\log\mathcal{N}(x;\mu,\Sigma)\right\}$
$=$ $-\frac{1}{2}\left[\Sigma^{-1}-\Sigma^{-1}\mathbb{E}_{p(x)}\left\{(x-\mu)(x-\mu)^{\top}\right\}\Sigma^{-1}\right].$
Setting this derivative equal to zero gives
$\Sigma^{-1}=\Sigma^{-1}\mathbb{E}_{p(x)}\left\{(x-\mu)(x-\mu)^{\top}\right\}\Sigma^{-1},$
which implies (11c). ∎
-B Proof of Theorem 1
Proof.
This proof is inspired by Section 6.2.2 in [45]. According to (16) and (26), we have
$\left.\frac{\partial^{2}J}{\partial\hat{x}_{t}^{2}}\right|_{v^{(i)}}=$ $\left.\frac{\partial^{2}}{\partial\hat{x}_{t}^{2}}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t},P_{t})}\left\{\ell(x_{t},y_{t})\right\}\right|_{v^{(i)}}+P_{t|t-1}^{-1}$ (52)
$=$ $\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(i)},P_{t}^{(i)})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial{x}_{t}^{2}}\right\}+P_{t|t-1}^{-1}$
$=$ $\left(P_{t}^{-1}\right)^{(i+1)}.$
According to (25) and (26), we have
$\left.\frac{\partial J}{\partial\hat{x}_{t}}\right|_{v^{(i)}}$ $=-\left(P_{t}^{-1}\right)^{(i+1)}\delta\hat{x}_{t}^{(i)},$ (53)
$\left.\frac{\partial J}{\partial P_{t}^{-1}}\right|_{v^{(i)}}$ $=-\frac{1}{2}\left(P_{t}\right)^{(i)}\delta P_{t}^{(i)}\left(P_{t}\right)^{(i)}.$
Combined with (52) and (53), the Tayor-series expansion can be expressed as
$J_{t}^{(i+1)}-J_{t}^{(i)}$
$\approx$ $-\frac{1}{2}\delta\hat{x}_{t}^{\top}\left(P_{t}^{-1}\right)^{(i+1)}\delta\hat{x}_{t}-\frac{1}{2}\mathrm{Tr}\left(P_{t}^{(i)}\delta P_{t}^{-1}P_{t}^{(i)}\delta P_{t}^{-1}\right)$
$=$ $-\frac{1}{2}\delta\hat{x}_{t}^{\top}\left(P_{t}^{-1}\right)^{(i+1)}\delta\hat{x}_{t}$
$-\frac{1}{2}\mathrm{vec}(\delta P_{t}^{-1})^{\top}\left(P_{t}^{(i)}\otimes P_{t}^{(i)}\right)\mathrm{vec}(\delta P_{t}^{-1})$
$\leq$ $0$
-C Proof of Corollary 1
Proof.
For linear Gaussian systems with output probability $p(y_{t}|x_{t})=\mathcal{N}(y_{t};Cx_{t},R)$, the one-step iteration of the covariance matrix in (26) is given by:
$\left(P_{t}^{-1}\right)^{(1)}=$ $P_{t|t-1}^{-1}+\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(0)},P_{t}^{(0)})}\left\{\frac{\partial^{2}\ell(x_{t},y_{t})}{\partial x_{t}^{2}}\right\},$ (54)
$=$ $P_{t|t-1}^{-1}+C^{\top}R^{-1}C.$
Comparing (54) with KF covariance update equation (21), we observe that $P_{t}^{(1)}=P_{t|t}$. Then, refering to (19) and using (54), the one-step iteration of the mean vector in (26) can be expressed as:
$\hat{x}^{(1)}_{t}=$ $\hat{x}^{(0)}_{t}-P_{t}^{(1)}\mathbb{E}_{\mathcal{N}(x_{t};\hat{x}_{t}^{(0)},P_{t}^{(0)})}\left\{\frac{\partial\ell(x_{t},y_{t})}{\partial x_{t}}\right\}$ (55)
$-P_{t}^{(1)}P_{t|t-1}^{-1}\left(\hat{x}_{t}^{(0)}-\hat{x}_{t|t-1}\right)$
$=$ $\hat{x}^{(0)}_{t}+P_{t|t}C^{\top}R^{-1}(y_{t}-C\hat{x}_{t}^{(0)})$
$-P_{t|t}\left(P_{t|t}^{-1}-C^{\top}R^{-1}C\right)\left(\hat{x}_{t}^{(0)}-\hat{x}_{t|t-1}\right)$
$=$ $\hat{x}_{t|t-1}+P_{t|t}C^{\top}R^{-1}\left(y_{t}-C\hat{x}_{t|t-1}\right).$
Furthermore, noting that:
$P_{t|t}C^{\top}R^{-1}$
$=$ $P_{t|t-1}C^{\top}R^{-1}$
$-P_{t|t-1}C^{\top}(R+CPC^{\top})^{-1}CP_{t|t-1}C^{\top}R^{-1}$
$=$ $P_{t|t-1}C^{\top}\big{(}I-(R+CP_{t|t-1}C^{\top})^{-1}(CP_{t|t-1}C^{\top}+R)$
$-(R+CPC^{\top})^{-1}R\big{)}R^{-1}$
$=$ $P_{t|t-1}C^{\top}(R+CP_{t|t-1}C^{\top})^{-1},$
we recognize that $P_{t|t}C^{\top}R^{-1}$ is the Kalman gain $K_{t}$. Therefore, (55) becomes:
$\hat{x}^{(1)}_{t}=\hat{x}_{t|t-1}+K_{t}\left(y_{t}-C\hat{x}_{t|t-1}\right),$
which is the standard KF mean update equation. Thus, the one-step iteration of the NANO filter in (26) yields the optimal estimate, regardless of the initialization $\hat{x}^{(0)}_{t}$ and $P^{(0)}_{t}$. ∎